# Javier Alejo, Antonio F. Galvao, Julian Martinez-Iriarte, and Gabriel Montes-Rojas

### v2

@@ 3 Identification: a control function approach
In this section, we propose sufficient conditions, together with a control function (CF) approach, to identify and estimate the parameters of interest in the model given in equations (1)–(2), in the presence of both endogeneity and general heteroskedasticity.
@@
Control Function Condition: $E(\varepsilon|D,X,Z)=E(\varepsilon|V)=r(V)$.
@@
$=D\alpha_{1}+X^{\prime}\alpha_{2}+g(D,X)r(V).$ (6)
@@
where $V:=\psi(D,Z)$ is an assumed identified control function from the first-stage. Now consider the following derivatives
$\frac{\partial f}{\partial D}$ $=\alpha_{1}+g_{D}r+gr_{V}\psi_{D}$
$\frac{\partial f}{\partial Z}$ $=gr_{V}\psi_{Z},$
@@
Polynomial Control Function: $r(V)$ is a polynomial of degree $k_{v}$ in $V$ that does not contain a constant. That is $r(V)=\sum_{j=1}^{k_{v}}\mu_{j}V^{j}$.
An important condition in Assumption 5 is the exclusion of a constant term in the function $r(V)$. Otherwise, we may allow for $r(V)$ to include a constant, but restrict $g(D,X)$ to not contain a linear term with $D$ and $X$. The intuition for these restrictions is simple. Suppose there is only one endogenous regression $D$ and the both the skedastic function and the CF contain a constant as: $g(D)=1+\delta D$ and $r(V)=1+V$. Then, $g(D)r(V)=1+V+\delta D+\delta DV$, and in this case, one is not able to identify $\alpha_{1}$ because of the term $\delta D$. This is also highlighted by Florens, Heckman, Meghir, and Vytlacil (2008).
Identification of the parameter of interest requires identification of $V$, which in turn requires identification of $h$, which we can achieved up-to-scale. See for instance, Romano and Wolf (2017) and Wooldridge (2012, ch. 8) for parametric specifications, and Jin, Su, and Xiao (2015) and Linton and Xiao (2019) for nonparametric specifications. The next assumption imposes a normalization on the second moment of the control function, and requires the skedastic function to be positive.
Assumption 6.
Either $h\equiv 1$ or $E[V^{2}|X,Z]=1$ and $h(X,Z)>0$ almost surely.
The next lemma considers the identification of $\alpha_{1}$. It shows that the parameter of interest is identified by a regression of $Y$ on transformations of $(D,X,V)$.
@@
$E[Y|D,X,Z]$ $=D\alpha_{1}+X^{\prime}\alpha_{2}+\left(1+\sum_{s=1}^{k_{g}}(\theta_{s}^{d}D^{s}+X^{\prime s}\theta_{s}^{x})\right)\left(\sum_{j=1}^{k_{v}}\mu_{j}V^{j}\right),$ (8)
@@ 4.1 An augmented control function estimator
We consider an estimator that consists of a two-step procedure. The overall procedure is as following. In the first-stage, one estimates the control function (CF), $V$. In the second-stage, one performs a simple OLS regression of the dependent variable $Y$ on the regressors $(D,X^{\prime})^{\prime}$ and the estimated CF, $\widehat{V}$, as well as on their polynomial interactions.
@@
Denote the preliminary estimator of $V$ denoted by $\widehat{V}$. The estimator is given by
@@
where $\widehat{L}[D_{i}|\widehat{W}_{i}]=\widehat{W}_{i}^{\prime}(\widehat{W}^{\prime}\widehat{W})^{-1}\widehat{W}^{\prime}D.$ Here $\widehat{W}_{i}$ denotes the vector $W_{i}$ with $V_{i}$ replaced by $\widehat{V}_{i}$.
To implement this strategy in practice, we first estimate the errors in the first-stage regression. To this end, we employ OLS in the first-stage and regress the endogenous variable $D$ on the instruments, $Z$, and exogenous variables $X$, to compute the residuals
$\widetilde{V}_{i}$ $=D_{i}-Z_{i}^{\prime}\widehat{\pi}_{1}+X_{i}^{\prime}\widehat{\pi}_{2}.$
@@
$\widehat{V}_{i}=\widetilde{V}_{i}/\widehat{h}_{i},$
@@
The coefficients of these models can be estimated using (possibly non-linear) OLS regressions with $\widetilde{V}^{2}_{i}$ as the dependent variable. The resulting estimator is $\widehat{h}_{i}=h(X_{i},Z_{i};\widehat{\gamma})$.
@@ 4.2 Asymptotic properties
From equation (2), $V$ is given by
@@
where $W(\phi)$ contains all the regressors except $D$, and by construction, $E[U|D,W]=0$. The following assumption is used in the proposition below to derive the influence function of $\widehat{\phi}$ and to derive the asymptotic properties of $\alpha_{1}$, the parameter of interest.
Assumption 7.
(i) The map $\gamma\mapsto h(x,z;\gamma)$ is twice differentiable with gradient denoted by $\nabla_{\gamma}h(x,z;\gamma)$; (ii) $E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}]$ has full rank; and (iii) the random vectors $Z_{i}h(X_{i},Z_{i})V_{i}$, $X_{i}h(X_{i},Z_{i})V_{i}$ and $\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}$ have finite second moments.
Assumption 7 imposes a standard regularity condition on the skedastic function determining the smoothness and existence of moments of the conditional heteroskedasticity. The next result establishes a linear representation for the first-step estimator.
@@ 5 Monte Carlo Simulations
We present results for different estimators: OLS, standard 2SLS (note that this is equivalent to a standard CF model where $\tilde{V}$ is included as an additional regressor in the structural equation), and the proposed CF with one interaction term CF1 $(\widehat{V},\widehat{V}D)$, as well as with two interaction terms CF2 $(\widehat{V},\widehat{V}D,\widehat{V}D^{2})$.
For the CF models we consider one type of skedastic function that matches the data generating process (18) given in Romano and Wolf (2017), as given in equation (10), $h(X,Z;\gamma)^{2}=|Z^{\prime}|\gamma_{1}+|X^{\prime}|\gamma_{2}=\gamma_{0}+|Z^{\prime}|\gamma_{1}$, such that $\hat{h}=\sqrt{h(X,Z;\hat{\gamma})^{2}}$. Let $\tilde{V}$ be the residuals of the first-stage regression of $D$ on $Z$ and a constant, then $h^{2}$ is estimated as the regression of $\tilde{V}^{2}$ on $|Z|$ and a constant.
@@
CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
@@
CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
@@
CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
@@
CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
@@ 6 Empirical application
Next, we estimate the effect of JTPA training on log earnings. Results for different models appear in Table 5 where only the coefficient of JTPA training is reported. Besides OLS and 2SLS, we also consider several control function methods. We present estimate results for different estimates of heteroskedasticity in the first-stage (Heter. first-stage). The first column of results, where Heter. first-stage $h=1$, does not consider a skedastic function in the first-stage and uses just the residuals $\tilde{V}$. The second column (Heter. first-stage: $h=V^{2}$) applies a first-stage skedastic function using a regression of $\tilde{V}^{2}$ on all covariates. The third column (Heter. first-stage: $h=ln\ V^{2}$) repeats a similar exercise but using a regression of $log\ \tilde{V}^{2}$ on all covariates. Regarding the CF model, the first row uses the estimated CF from the first-stage ($\tilde{V}$). The second row considers the residual and its squared value ($\widehat{V}+\widehat{V^{2}}$). The third row adds the interaction with JTPA training to accommodate a given skedastic function in the structural equation ($\widehat{V}+\widehat{V}D$). The fourth row adds the interaction and the square ($\widehat{V}+\widehat{V}D+\widehat{V}^{2}$). Finally, the last row considers the case $\widehat{V}+\widehat{V}D+\widehat{V}^{2}+\widehat{V}^{2}D$. Standard errors are computed by bootstrap using 200 replications.
@@
The empirical results in Table 5 show that the OLS point estimate is 0.266, while the 2SLS estimate is 0.115. This shows a substantial decrease in effect of training on earning, which indicates that the 2SLS correction with respect to OLS might be too large. Next, we consider the control function estimates. Note that the result for the first row ($\widehat{V}$) and first column ($h=1$) coincides with 2SLS estimate. This is because in this case there is no heteroskedasticity in the first-stage, and the second-stage does not consider any interactions. The presence of heteroskedasticity documented before indicates that models using interactions as control function might be more appropriate. Thus, columns $h=V^{2}$ and $h=\ln(V^{2})$ are important. For the case where $\widehat{V}+\widehat{V}D$, the point estimate ranges from 0.189 to 0.234, depending on the first-stage correction (and discarding the model with no correction for heteroskedasticity in the first-stage). For the case $\widehat{V}+\widehat{V}D+\widehat{V}^{2}$ estimates range from 0.20 to 0.227, while for $\widehat{V}+\widehat{V}D+\widehat{V}^{2}+\widehat{V}^{2}D$ it ranges from 0.180 to 0.235. These cases show that results are robust to the inclusion of polynomial and interaction terms. Overall, the CF results show empirical evidence that the job training is effective in increasing earning, but estimates are more modest than those from simple OLS regressions.
@@ 7 Conclusion
This paper shows that the estimation of linear models with endogeneity using the popular two-stages least squares (2SLS) method may not be estimating the parameter of interest. In particular, in the presence of heteroskedasticity, something common when correcting the variance-covariance matrix and for doing correct inference, the 2SLS is inconsistent. The empirical literature has not paid specific attention to this issue, probably because the interest lied in estimating the LATE rather than the ATE, as popularized by the treatment effects agenda.
As an alternative we provide a simple to implement augmented control function estimator. The control function (CF) approach provides an intuitive modeling strategy to handle endogeneity (see Wooldridge (2015) for a review) and lies in the center of general approaches to this particular set-up (see Newey, Powell, and Vella (1999) and Florens, Heckman, Meghir, and Vytlacil (2008)). We show that this can be adapted to the least-squares linear estimation methods to correct for the 2SLS bias. In fact, the proposed solution is to augment the standard CF model (which corresponds to 2SLS) with interactions of the structural equation variables with the residuals from the first-stage.
An important future research is the extension of the methods developed in this paper to study panel data models.
@@ A.3 Proof of Lemma 1
Recall that the model is
@@
This is the general form of the bias, without taking into account the functional form relating $D$ to $Z$ and $X$. Note right away in (A.3), that if $g(D,X)\equiv 1$, then there is no bias, since $E[\overline{Z}\,\overline{\varepsilon}]=0$ by the validity of $Z$.
Apply a linear projection to the first-stage:
$L_{[D|X]}(X)$ $=L_{[Z^{\prime}\pi_{1}|X]}(X)+X^{\prime}\pi_{2}+L_{[h(X,Z)V|X]}(X)$
Therefore, the residualized version is
$\overline{D}$ $=\overline{Z}^{\prime}\pi_{1}+\overline{hV}.$ (A.4)
where $h:=h(X,Z)$ for notational simplicity. Now we plug-in (A.4) into the bias part of (A.3). First note that
$E[\overline{Z}\,\overline{D}]$ $=E[\overline{Z}\,\overline{Z}^{\prime}]\pi_{1}+E[\overline{Z}\,\overline{hV}]$
@@
because, by Assumption 1, $E[V|X,Z]=0$, which implies $E[\overline{Z}\,\overline{hV}]=0.$ So that
@@
$\overline{\alpha}_{1}=\alpha_{1}+\Sigma_{h}^{-1}\pi_{1}^{\prime}E[\overline{Z}\,\overline{g\varepsilon}].$
Now, since $g:=g(D,X)$, and $D=Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}+h(X,Z)V,$ the final expression for the asymptotic bias is
bias $=\Sigma_{h}^{-1}\pi_{1}^{\prime}E[\overline{Z}g(Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}+h(X,Z)V,X)\varepsilon].$

### v3

@@ 3 Identification: a control function approach
In this section, we propose sufficient conditions, together with a control function approach, to identify and estimate the parameters of interest in the model given in equations (1)–(2), in the presence of both endogeneity and general heteroskedasticity.
@@
Control Function Condition: there exists $\tilde{V}$ which satisfies $E[\varepsilon|D,X,Z]=E[\varepsilon|\tilde{V}]=r(\tilde{V})$.
@@
$=D\alpha_{1}+X^{\prime}\alpha_{2}+g(D,X)r(\tilde{V}).$ (4)
@@
where $\tilde{V}:=\psi(D,Z)$ is an assumed identified control function from the first-stage. Now consider the following derivatives
$\frac{\partial f}{\partial D}$ $=\alpha_{1}+g_{D}r+gr_{\tilde{V}}\psi_{D}$
$\frac{\partial f}{\partial Z}$ $=gr_{\tilde{V}}\psi_{Z},$
@@
Polynomial Control Function: $r(\tilde{V})$ is a polynomial of degree $k_{v}$ in $\tilde{V}$ that does not contain a constant. That is $r(\tilde{V})=\sum_{j=1}^{k_{v}}\mu_{j}\tilde{V}^{j}$.
An important condition in Assumption 5 is the exclusion of a constant term in the function $r(\tilde{V})$. Otherwise, we may allow for $r(\tilde{V})$ to include a constant, but restrict $g(D,X)$ to not contain a linear term with $D$ and $X$. The intuition for these restrictions is simple. Suppose there is only one endogenous regression $D$ and the both the skedastic function and the CF contain a constant as: $g(D)=1+\delta D$ and $r(\tilde{V})=1+\tilde{V}$. Then, $g(D)r(\tilde{V})=1+\tilde{V}+\delta D+\delta D\tilde{V}$, and in this case, one is not able to identify $\alpha_{1}$ because of the term $\delta D$. This is also highlighted by Florens, Heckman, Meghir, and Vytlacil (2008).
The next lemma considers the identification of $\alpha_{1}$. It shows that the parameter of interest is identified by a regression of $Y$ on transformations of $(D,X,\tilde{V})$.
@@
$E[Y|D,X,Z]$ $=D\alpha_{1}+X^{\prime}\alpha_{2}+\left(1+\sum_{s=1}^{k_{g}}(\theta_{s}^{d}D^{s}+X^{\prime s}\theta_{s}^{x})\right)\left(\sum_{j=1}^{k_{v}}\mu_{j}\tilde{V}^{j}\right),$ (6)
@@ 4.1 An augmented control function estimator
We consider an estimator that consists of a two-step procedure. The overall procedure is as following. In the first-stage, one estimates the control function (CF), $\tilde{V}$. In the second-stage, one performs a simple OLS regression of the dependent variable $Y$ on the regressors $(D,X^{\prime})^{\prime}$ and the estimated CF, $\widehat{V}$, as well as on their polynomial interactions.
@@
Denote the preliminary estimator of $\tilde{V}$ denoted by $\widehat{V}$. The estimator is given by
@@
where $\widehat{L}[D_{i}|\widehat{W}_{i}]=\widehat{W}_{i}^{\prime}(\widehat{W}^{\prime}\widehat{W})^{-1}\widehat{W}^{\prime}D.$ Here $\widehat{W}_{i}$ denotes the vector $W_{i}$ with $\tilde{V}_{i}$ replaced by $\widehat{V}_{i}$.
To implement this strategy in practice, we first estimate the errors in the first-stage regression assuming a linear heteroskedastic model.
Assumption 6.
Assume $D=m(X,Z,V)=Z^{\prime}\pi_{1}+X^{\prime}\pi 2+h(X,Z)V$, with $E[V^{2}|X,Z]=1$ and $|h(X,Z)|>0$ almost surely.
The CF would be then the homoskedastic transformation of the first-stage residual, i.e. $\tilde{V}=V$ as in Assumption 6. To this end, we employ OLS in the first-stage and regress the endogenous variable $D$ on the instruments, $Z$, and exogenous variables $X$, to compute the residuals
$\check{V}_{i}$ $=D_{i}-Z_{i}^{\prime}\widehat{\pi}_{1}+X_{i}^{\prime}\widehat{\pi}_{2}.$
@@
$\widehat{V}_{i}=\check{V}_{i}/\widehat{h}_{i},$
@@
The following assumption imposes a standard regularity condition on the skedastic function determining the smoothness and existence of moments of the conditional heteroskedasticity.
Assumption 7.
Assume that $h(X,Z)=h(X,Z;\gamma)$ a parametric function with parameter $\gamma$. (i) The map $\gamma\mapsto h(x,z;\gamma)$ is twice differentiable with gradient denoted by $\nabla_{\gamma}h(x,z;\gamma)$; (ii) $E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}]$ has full rank; and (iii) the random vectors $Z_{i}h(X_{i},Z_{i})V_{i}$, $X_{i}h(X_{i},Z_{i})V_{i}$ and $\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}$ have finite second moments.
@@
The coefficients of these models can be estimated using (possibly non-linear) OLS regressions with $\check{V}^{2}_{i}$ as the dependent variable. The resulting estimator is $\widehat{h}_{i}=h(X_{i},Z_{i};\widehat{\gamma})$.
@@ 4.2 Asymptotic properties
From Assumptions 6 and 7, $V$ is given by
@@
where $W(\phi)$ contains all the regressors except $D$, and by construction, $E[U|D,W]=0$.
The next result derive the influence function of $\widehat{\phi}$ and to derive the asymptotic properties of $\alpha_{1}$, the parameter of interest.
@@ 5 Monte Carlo Simulations
We present results for different estimators: OLS, standard 2SLS (note that this is equivalent to a standard CF model where $\check{V}$ is included as an additional regressor in the structural equation), and the proposed CF with one interaction term CF1 $(\widehat{V},\widehat{V}D)$, as well as with two interaction terms CF2 $(\widehat{V},\widehat{V}D,\widehat{V}D^{2})$.
For the CF models we consider one type of skedastic function that matches the data generating process (16) given in Romano and Wolf (2017), as given in equation (8), $h(X,Z;\gamma)^{2}=|Z^{\prime}|\gamma_{1}+|X^{\prime}|\gamma_{2}=\gamma_{0}+|Z^{\prime}|\gamma_{1}$, such that $\hat{h}=\sqrt{h(X,Z;\hat{\gamma})^{2}}$. Let $\check{V}$ be the residuals of the first-stage regression of $D$ on $Z$ and a constant, then $h^{2}$ is estimated as the regression of $\check{V}^{2}$ on $|Z|$ and a constant.
@@
OLS 2SLS CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
@@
OLS 2SLS CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
@@
OLS 2SLS CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
@@
OLS 2SLS CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
@@ 6 Empirical application
Next, we estimate the effect of JTPA training on log earnings. Results for different models appear in Table 5 where only the coefficient of JTPA training is reported. Besides OLS and 2SLS, we also consider several control function methods. We present estimate results for different estimates of heteroskedasticity in the first-stage (Heter. first-stage). The first column of results, where Heter. first-stage $h=1$, does not consider a skedastic function in the first-stage and uses just the residuals $\check{V}$. The second column (Heter. first-stage: $h=V^{2}$) applies a first-stage skedastic function using a regression of $\check{V}^{2}$ on all covariates. The third column (Heter. first-stage: $h=ln\ V^{2}$) repeats a similar exercise but using a regression of $log\ \check{V}^{2}$ on all covariates. Regarding the CF model, the first row uses the estimated CF from the first-stage ($\check{V}$). The second row considers the residual and its squared value $(\widehat{V},\widehat{V}^{2})$. The third row adds the interaction with JTPA training to accommodate a given skedastic function in the structural equation $(\widehat{V},\widehat{V}D)$. The fourth row adds the interaction and the square ($\widehat{V},\widehat{V}D,\widehat{V}^{2}$). Finally, the last row considers the case $(\widehat{V},\widehat{V}D,\widehat{V}^{2},\widehat{V}^{2}D)$. Standard errors are computed by bootstrap using 200 replications.
@@
The empirical results in Table 5 show that the OLS point estimate is 0.266, while the 2SLS estimate is 0.115. Next, we consider the CF estimates. Note that the result for the first row ($\widehat{V}$) and first column ($h=1$) coincides with 2SLS estimate. This is because in this case there is no heteroskedasticity in the first-stage, and the second-stage does not consider any interactions. The presence of heteroskedasticity documented before indicates that models using interactions as CF might be more appropriate. Thus, columns $h=V^{2}$ and $h=\ln(V^{2})$ are important. For the case where $(\widehat{V},\widehat{V}D)$, the point estimate ranges from 0.189 to 0.234, depending on the first-stage correction (and discarding the model with no correction for heteroskedasticity in the first-stage). For the case $(\widehat{V},\widehat{V}D,\widehat{V}^{2})$ estimates range from 0.20 to 0.227, while for $(\widehat{V},\widehat{V}D,\widehat{V}^{2},\widehat{V}^{2}D)$ it ranges from 0.180 to 0.235. These cases show that results are robust to the inclusion of polynomial and interaction terms. Overall, the CF results show empirical evidence that the job training is effective in increasing earning, but estimates are more modest than those from simple OLS regressions.
@@ 7 Conclusion
This paper shows that the estimation of linear models with endogeneous heteroskedasticity using the popular 2SLS method may not be estimating the parameter of interest. As an alternative we provide a simple to implement augmented control function estimator. The CF approach provides an intuitive modeling strategy to handle endogeneity and lies in the center of general approaches to this particular set-up. We show that this can be adapted to the least-squares linear estimation methods to correct for the 2SLS bias. In fact, the proposed solution is to augment the standard CF model (which corresponds to 2SLS) with interactions of the structural equation variables with the residuals from the first-stage.
@@ A.3 Proof of Lemma 1
Recall that the model in (1)–(2) is
@@
Let $V_{D}:=D-L_{[D|Z,X]}(Z,X)$ be the residuals from the linear projection of $D$ on $(Z,X)$, with $L_{[D|Z,X]}(Z,X):=Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}$. Thus we can write
$L_{[D|X]}(X)$ $=L_{[Z^{\prime}\pi_{1}|X]}(X)+X^{\prime}\pi_{2}+\underbrace{L_{[V_{D}|X]}(X)}_{=0.}$
but since $L_{[V_{D}|Z,X]}(Z,X)=0$, then $L_{[V_{D}|X]}(X)=0$. Therefore, the residualized version of $D$ with respect to $X$ is
$\underbrace{\overline{D}}_{D-L_{[D|X]}(X)}$ $=\overline{Z}^{\prime}\pi_{1}+V_{D}$ (A.4)
Now we plug-in (A.4) into the bias part of (A.3). First note that
$E[\overline{Z}\,\overline{D}]$ $=E[\overline{Z}\,\overline{Z}^{\prime}]\pi_{1}+E[\overline{Z}V_{D}]$
@@
because $V_{D}$ is orthogonal to any linear function of both $X$ and $Z$, such as $\overline{Z}^{\prime}:=Z^{\prime}-L_{[Z^{\prime}|X]}(X)$. So that
@@
$\overline{\alpha}_{1}=\alpha_{1}+\Sigma_{Z}^{-1}\pi_{1}^{\prime}E[\overline{Z}\,\overline{g\varepsilon}].$
Now, since $g:=g(D,X)$, and $D=m(X,Z,V),$ the final expression for the asymptotic bias is
bias $=\Sigma_{Z}^{-1}\pi_{1}^{\prime}E[\overline{Z}g(m(X,Z,V),X)\varepsilon].$